OpenStudy (anonymous):

Solve for pi: C = 2pir

OpenStudy (anonymous):

\[c=2piR\]

OpenStudy (debbieg):

Well, you need to get the \(\pi\) "alone". Currently it is multiplied by 2r. So what do you need to do, to "undo" that multiplication, and the the \(\pi\) all by itself?

OpenStudy (anonymous):

but how would i get pi alone

OpenStudy (anonymous):

so it would be B

OpenStudy (debbieg):

If \(a=bc\) then \(\dfrac{a}{c}=b\) "undo" multiplication by dividing both sides by the necessary factor.

OpenStudy (anonymous):

oh

OpenStudy (debbieg):

I don't have your answer choices, but that's ok - I'd rather you understand how to do the problem. then YOU will KNOW which answer choice is correct.

OpenStudy (anonymous):

ok thxs

OpenStudy (debbieg):

If I have \(7k=3az\) and I want to get z alone, I divide both sides by 3a: \(\dfrac{7k}{3a}=\dfrac{3az}{3a}\) \(\dfrac{7k}{3a}=z\)

OpenStudy (anonymous):

ok